# ParaICL: Towards Robust Parallel In-Context Learning

### v1

@@ Abstract
Large language models (LLMs) have become the norm in natural language processing (NLP), excelling in few-shot in-context learning (ICL) with their remarkable abilities. Nonetheless, the success of ICL largely hinges on the choice of few-shot demonstration examples, making the selection process increasingly crucial. Existing methods have delved into optimizing the quantity and semantic similarity of these examples to improve ICL performances. However, our preliminary experiments indicate that the effectiveness of ICL is limited by the length of the input context. Moreover, varying combinations of few-shot demonstration examples can significantly boost accuracy across different test samples. To address this, we propose a novel method named parallel in-context learning (ParaICL) that effectively utilizes all demonstration examples without exceeding the manageable input context length. ParaICL employs parallel batching to distribute demonstration examples into different batches according to the semantic similarities of the questions in the demonstrations to the test question. It then computes normalized batch semantic scores for each batch. A weighted average semantic objective, constrained by adaptive plausibility, is applied to select the most appropriate tokens. Through extensive experiments, we validate the effectiveness of ParaICL and conduct ablation studies to underscore its design rationale. We further demonstrate that ParaICL can seamlessly integrate with existing methods  [^1] We will make our code publicly available..
@@ 1 Introduction
In recent years, scaling up the parameters of generative language models has significantly enhanced their language generation capabilities (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023). Large language models (LLM) have demonstrated their adeptness across a wide range of tasks via few-shot in-context learning (ICL) (Cheng et al., 2023; Zhao et al., 2023; Li et al., 2024). In few-shot ICL, models are expected to generate outputs directly based on a sequence of given examples without the need for any parameter modifications, making it the most efficient method for adapting to new tasks.
However, the effectiveness of ICL is notably influenced by the few-shot demonstration examples used (Chen et al., 2023). Various methods have been developed to select the most effective few-shot demonstration examples for ICL. Hao et al. (2022) demonstrated that scaling up the number of demonstration examples can improve the ICL performance. Rubin et al. (2022) and Liu et al. (2022) proposed to utilize the most relevant examples for each test sample during inference. Nevertheless, existing few-shot ICL methods mostly focus on either increasing the number of demonstration examples or selecting a few that are semantically similar to the test samples to improve performance.
@@
To combine the best of both worlds, we introduce parallel in-context learning (ParaICL). ParaICL effectively utilizes the maximum number of demonstration examples without extending the input context length, thus avoiding the potential reduction in model performance due to larger context sizes. Given a set of question–answer pairs as demonstration examples, ParaICL first assigns them into various batches based on the semantic similarity between the demonstrations’ questions and the sample test question. Consequently, each batch maintains a controlled context length while utilizing all demonstration examples simultaneously. These batches are then processed by a causal model in parallel to obtain the next token distribution for each batch. Afterward, a weighted average of these distributions is calculated, considering the semantic relation of each batch to the test question. The final step involves selecting the token with the highest weighted average probability for continued generation.
We conduct extensive experiments across a range of reasoning, natural language inference, and coding tasks to validate the effectiveness of ParaICL. We further demonstrate that ParaICL is compatible with both open-source and black-box causal language models. Our study includes abundant ablation studies and analyses to justify the design of ParaICL and demonstrate how it can be integrated with other ICL methods. In summary, our main contributions are the following: (1) We introduce parallel in-context learning (ParaICL), a simple but effective method that leverages all available demonstration examples while maintaining the input context length manageable. (2) We conduct thorough experiments to prove the effectiveness of our method, along with ablation studies to justify its design. (3) We illustrate how our method can enhance and work in conjunction with other methods.
@@ 2 Related Work
In-context learning (ICL) has surged as a transformative approach in the natural language processing (NLP) domain, primarily enhancing the adaptability and efficiency of causal large language models (LLMs) (Brown et al., 2020). The significance of ICL is evident across various applications, including knowledge grounding (Zhao et al., 2023; Li et al., 2024), code generation (Li et al., 2023b), and other industrial applications (Cheng et al., 2023; Chen et al., 2024). Despite its promise, challenges such as sensitivity to the prompts and context window length limitations, highlight areas for further research and development in making ICL more robust and versatile.
As aforementioned, various studies have underscored that ICL exhibits significant sensitivity to the quality of prompts, particularly with regard to the demonstration examples provided. Chen et al. (2023) discovered that increasing the number of demonstration examples only leads to slight improvements. Liu et al. (2022) proposed to select examples that have the highest semantic similarity to the test question. Contrarily, Levy et al. (2023) found that leveraging a diverse set of demonstration examples could improve in-context compositional generalization. Qin et al. (2023) proposed iterative demonstration selection, which considers both the diversity and similarity dimensions of ICL demonstration selection for LLMs. Additionally, Chia et al. (2023) explored the potential of contrastive examples in improving the reasoning capabilities of causal language models. Nevertheless, these methods primarily focus on choosing a subset of demonstration examples from a pool of candidates based on certain perspectives. They fall short in striking a delicate balance between the quantity and the quality of the demonstration examples.
Another line of research is dedicated to utilizing extended context in ICL. Ratner et al. (2023) proposed parallel context window (PCW), a method that alleviates the context window limitations for any off-the-shelf LLMs without necessitating additional training. However, PCW overlooked the semantic connections between the demonstration examples and the test samples. As evidenced by Chen et al. (2023), merely increasing the quantity of demonstration examples does not correlate with a substantial improvement in performance. In fact, this strategy might detract from overall performance due to the inclusion of incorrect or misleading content within some of the demonstration examples.
@@ 3 Methodology
We first formulate the problem setting for few-shot in-context learning (ICL) in Section 3.1. Following this, in Section 3.2, we introduce parallel in-context learning (ParaICL), a novel method that efficiently processes demonstration examples in batches and effectively utilizes a novel parallel semantic decoding strategy for generation. Subsequently in Section 3.3, we introduce the weighted average semantic objective alongside the adaptive plausibility constraint. Finally we define the parallel semantic decoding method, which employs the weighted average semantic objective subject to the plausibility constraint. A demonstration of ParaICL can be found in Figure 3.
@@ 3.3 Parallel Semantic Decoding
We first define a generative language model denoted as $f_{lm}(\cdot)$. For given inputs $\mathrm{\textbf{u}}_{prev}^{i}$, the model generates a continuation $\mathrm{\textbf{u}}_{cont}^{i}$ according to the formula:
@@
As demonstration examples have varying semantic contributions to the test sample, we propose the weighted average semantic (WAS) objective. The WAS objective for test sample $\hat{x}_{i}$ is defined as: {dmath} L_WAS(u_cont^i, O^i, U^i) = ∑_z=1^so_z^i ⋅f_lm(u_cont^i—u_z^i). The WAS objective rewards batches with demonstration examples that exhibit greater semantic similarity to the test question, while simultaneously leveraging all examples to enrich the generation process comprehensively. However, certain batches may contain noise that could adversely affect the performance. To address this problem, we adopt the adaptive plausibility constraint from the contrastive decoding method (Li et al., 2023a).
@@
$\mathcal{V}_{head}(u_{<j}^{i})=\{u_{j}^{i}\in\mathcal{V}:f_{lm}(u_{j}^{i}|u_{<j}^{i},\mathrm{\textbf{u}}_{1}^{i})\geq\alpha\max_{w^{i}}f_{lm}(w^{i}|u_{<j}^{i},\mathrm{\textbf{u}}_{1}^{i})\},$ (5)
@@
$\max_{\mathrm{\textbf{u}}_{cont}^{i}}{\mathcal{L}_{\mathrm{WAS}}(\mathrm{\textbf{u}}_{cont}^{i},\mathcal{O}^{i},\mathcal{U}^{i})}\quad\text{subject to }u_{j}^{i}\in\mathcal{V}_{head}(u_{<j}^{i}),\forall u_{j}^{i}\in\mathrm{\textbf{u}}_{cont}^{i}.$ (6)
@@
$\mathcal{L}_{\mathrm{WAS}}(\mathrm{\textbf{u}}_{cont}^{i},\mathcal{O}^{i},\mathcal{U}^{i})=\sum_{z=1}^{s}{o_{z}^{i}\cdot f_{lm}{(\mathrm{\textbf{u}}_{cont}^{i}|\mathrm{\textbf{u}}_{z}^{i})}}=\prod_{j=1}^{q^{i}}{\text{WAS-score}(u_{j}^{i},u_{<j}^{i},\mathcal{O}^{i},\mathcal{U}^{i})},$ (7)
@@
[…]
@@ 4 Experiments
4.1 Datasets
4.2 Baselines
To provide a more comprehensive overview of where our framework stands, we use the following baselines: (1) Standard few-shot (Standard) (Brown et al., 2020): Directly generating the results based on few-shot demonstration examples provided in sequence within the input prompt. (2) Semantically sorted few-shot (Sorted) (Chen et al., 2023): Utilizing the same few-shot demonstration examples, this approach organizes the examples by the semantic similarity between each example’s question and the test question. Sorted+ indicates examples sorted in ascending similarity and Sorted- in descending similarity. (3) Parallel context window (PCW) (Ratner et al., 2023): Carving a long context into batches, PCW restricts the attention mechanism to apply within each batch, and re-use the positional embeddings across the batches. To ensure a fair comparison, we maintain the same number of batches as used in ParaICL.
4.3 Experiment Setup for ParaICL
When setting up experiments for ParaICL, we maintain consistency with the demonstration examples used in baseline methods to ensure a fair comparison. To demonstrate the effectiveness of ParaICL, we conduct experiments with varying numbers of total demonstration examples, specifically 3, 9, and 15, while keeping the count of parallel batches at three. Consequently, this results in 1, 3, and 5 demonstration examples per batch across the three settings. The impact of batch numbers will be elaborated in Section 5.1. Each experimental run is carried out with three random seeds for the selection of demonstration examples, and the results are averaged for reporting. Our main experiments are conducted using open-source models, including Llama2-7B-Chat and Mistral-7B-Instruct-v0.2  [^3] Due to cost concern, the validation of ParaICL’s effectiveness on close-source models is conducted on a more compact dataset in Section 5.4.. We utilize supervised SimCSE with BERT base to compute sentence embeddings.
@@ 4.4 Experiment Results
We present the experimental results of all datasets from Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 in Table 1. ParaICL consistently outperforms baseline methods on all datasets, demonstrating the effectiveness of our method. On reasoning tasks, the average improvements with Llama-2-7B-Chat are 1.2%, 2.7%, and 2.0% for 3-shot, 9-shot, and 15-shot settings, respectively. Mistral-7B-Instruct-v0.2 shows improvements of 1.0%, 2.9%, and 3.0% for the same settings, demonstrating ParaICL’s adaptability across different large language models. On commonsense natural language inference (NLI) and coding tasks, ParaICL presents an average performance boost of 1.8% and 2.0% for Llama-2-7B-Chat, and 1.9% and 2.0% for Mistral-7B-Instruct-v0.2, respectively. Moreover, ParaICL consistently surpasses PCW in all settings and on virtually all datasets, underscoring the advantages of utilizing semantic information in our approach. An observation is made regarding the performance similarity between the Sorted+ and Sorted- methods. Sorted+ sometimes outperforms Sorted- and vice versa. This finding is consistent with prior research by Levy et al. (2023) and Liu et al. (2022). With ParaICL, this issue is lessened due to the reduced number of demonstration examples per batch. In ParaICL, we follow the common practices (Liu et al., 2022) to arrange the demonstrations in ascending order.

### v2

# ParaICL: Towards Parallel In-Context Learning

@@ Abstract
Large language models (LLMs) have become the norm in natural language processing (NLP), excelling in few-shot in-context learning (ICL) with their remarkable abilities. Nonetheless, the success of ICL largely hinges on the choice of few-shot demonstration examples, making the selection process increasingly crucial. Existing methods have delved into optimizing the quantity and semantic similarity of these examples to improve ICL performances. However, our preliminary experiments indicate that the effectiveness of ICL is limited by the length of the input context. Moreover, varying combinations of few-shot demonstration examples can significantly boost accuracy across different test samples. To address this, we propose a novel method named parallel in-context learning (ParaICL) that effectively utilizes all demonstration examples without exceeding the manageable input context length. ParaICL employs parallel batching to distribute demonstration examples into different batches according to the semantic similarities of the questions in the demonstrations to the test question. It then computes normalized batch semantic scores for each batch. A weighted average semantic objective, constrained by adaptive plausibility, is applied to select the most appropriate tokens. Through extensive experiments, we validate the effectiveness of ParaICL and conduct ablation studies to underscore its design rationale. We further demonstrate that ParaICL can seamlessly integrate with existing methods. Our code is available at https://github.com/xingxuanli/paraicl.git.
ParaICL: Towards Parallel In-Context Learning
Xingxuan Li1,2thanks: Work done while at DAMO Academy, Alibaba Group. Xingxuan Li was under the Joint Ph.D. Program between Alibaba and Nanyang Technological University.  Xuan-Phi Nguyen311footnotemark: 1  Shafiq Joty2,3  Lidong Bing [^1] footnotemark: 1thanks: Corresponding author. 1Shanda AI Research Institute 2Nanyang Technological University 3Salesforce Research {xingxuan.li, lidong.bing}@shanda.com {nguyenxu002, srjoty}@ntu.edu.sg
@@ 1 Introduction
In recent years, scaling up the parameters of generative language models has significantly enhanced their language generation capabilities (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023). Large language models (LLM) have demonstrated their adeptness across a wide range of tasks via few-shot in-context learning (ICL) (Cheng et al., 2023; Zhao et al., 2023; Li et al., 2024). In few-shot ICL, models are expected to generate outputs directly based on a sequence of given examples without any parameter modifications, making it the most efficient method for adapting to new tasks.
However, the effectiveness of ICL is notably influenced by the few-shot demonstration examples used (Chen et al., 2023). Various methods have been developed to select the most effective few-shot demonstration examples for ICL. Hao et al. (2022a) demonstrated that scaling up the number of demonstration examples can improve the ICL performance. Rubin et al. (2022) and Liu et al. (2022) proposed to utilize the most relevant examples for each test sample during inference. Nevertheless, existing few-shot ICL methods mostly focus on either increasing the number of demonstration examples or selecting a few that are similar to the test samples to improve performance.
@@
To combine the best of both worlds, we introduce parallel in-context learning (ParaICL). ParaICL effectively utilizes the maximum number of demonstration examples without extending the input context length, thus avoiding the potential reduction in model performance due to larger context sizes. Given a set of question–answer pairs as demonstration examples, ParaICL first assigns them into various batches based on the semantic similarity between the demonstrations’ questions and the sample test question. Consequently, each batch maintains a controlled context length while utilizing all demonstration examples. These batches are then processed by a causal model in parallel to obtain the next token distribution for each batch. Afterward, a weighted average of these distributions is calculated, considering the semantic relation of each batch to the test question. The final step involves selecting the token with the highest weighted average probability for continued generation.
We conduct extensive experiments across a range of reasoning, natural language inference, and coding tasks to validate the effectiveness of ParaICL. We further demonstrate that ParaICL is compatible with both open- and closed-source causal language models. Our study includes abundant ablation studies and analyses to justify the design of ParaICL and demonstrate how it can be integrated with other ICL methods. In summary, our main contributions are the following: (1) We introduce ParaICL, a simple but effective method that leverages all available demonstration examples while maintaining the input context length manageable. (2) We conduct thorough experiments to prove the effectiveness of our method, along with ablation studies to justify its design. (3) We illustrate how our method can enhance and work in conjunction with other methods.
@@ 2 Related Work
ICL has surged as a transformative approach in the NLP domain, with its significance evident across various applications, including knowledge grounding (Zhao et al., 2023; Li et al., 2024), code generation (Li et al., 2023b), and other industrial applications (Cheng et al., 2023; Chen et al., 2024). Despite its promise, challenges such as sensitivity to the prompts and context window length limitations, highlight areas for further research and development in making ICL more robust and versatile.
As aforementioned, various studies have underscored that ICL exhibits significant sensitivity to the quality of prompts, particularly concerning the demonstration examples provided (Li and Qiu, 2023a; Gupta et al., 2023; Raventos et al., 2023). Chen et al. (2023) discovered that increasing the number of examples only leads to slight improvements. Liu et al. (2022) proposed to select examples that have the highest semantic similarity to the test question. Contrarily, Levy et al. (2023) found that leveraging a diverse set of demonstration examples could improve in-context compositional generalization. Qin et al. (2023) proposed iterative demonstration selection, which considers both the diversity and similarity dimensions of ICL demonstration selection for LLMs. Additionally, Chia et al. (2023) explored the potential of contrastive examples in improving the reasoning capabilities of causal language models. Nevertheless, these methods primarily focus on choosing a subset of demonstration examples from a pool of candidates based on certain perspectives. Furthermore, these methods necessitate a substantial number of candidates to efficiently select the best examples, which constrains their applicability across all situations.
Another line of research is dedicated to utilizing extended context in ICL. Ratner et al. (2023) proposed parallel context window (PCW), a method that alleviates the context window limitations for any open-source LLMs without necessitating additional training. However, PCW overlooked the semantic connections between the demonstration examples and the test samples. As evidenced by Chen et al. (2023), merely increasing the quantity of demonstration examples does not correlate with a substantial improvement in performance. In fact, this strategy might detract from overall performance due to the inclusion of incorrect or misleading content within some of the examples.
@@ 3 Methodology
We first formulate the problem setting for few-shot ICL in Section 3.1. Following this, in Section 3.2, we introduce ParaICL, a novel method that efficiently processes demonstration examples in batches and effectively utilizes a novel parallel semantic decoding strategy for generation. Subsequently in Section 3.3, we introduce the weighted average semantic objective alongside the adaptive plausibility constraint. Finally we define the parallel semantic decoding method, which employs the weighted average semantic objective subject to the plausibility constraint. A demonstration of ParaICL can be found in Figure 3.
@@ 3.3 Parallel Semantic Decoding
We first define a generative language model $f_{lm}(\cdot)$. For given inputs $\mathrm{\textbf{u}}_{prev}^{i}$, the model generates a continuation $\mathrm{\textbf{u}}_{cont}^{i}$ according to the formula:
@@
As demonstration examples have varying semantic contributions to the test sample, we propose the weighted average semantic (WAS) objective. The WAS objective for test sample $\hat{x}_{i}$ is defined as: {dmath} L_WAS(u_cont^i, O^i, U^i) = ∑_z=1^so_z^i ⋅f_lm(u_cont^i|u_z^i). The WAS objective rewards batches with demonstration examples that exhibit greater semantic similarity to the test question, while simultaneously leveraging all examples to enrich the generation process comprehensively. This advantage makes ParaICL superior to PCW, which arranges examples randomly. However, certain batches may contain noise that could adversely affect the performance. To address this problem, we adopt the adaptive plausibility constraint from the contrastive decoding method (Li et al., 2023a).
@@
$\begin{split}&\mathcal{V}_{head}(u_{<j}^{i})\\ &=\{u_{j}^{i}\in\mathcal{V}:f_{lm}(u_{j}^{i}|u_{<j}^{i},\mathrm{\textbf{u}}_{1}^{i})\\ &\quad\quad\geq\alpha\max_{w^{i}}f_{lm}(w^{i}|u_{<j}^{i},\mathrm{\textbf{u}}_{1}^{i})\},\end{split}$ (5)
@@
$\begin{split}&\max_{\mathrm{\textbf{u}}_{cont}^{i}}{\mathcal{L}_{\mathrm{WAS}}(\mathrm{\textbf{u}}_{cont}^{i},\mathcal{O}^{i},\mathcal{U}^{i})}\\ &\text{subject to }u_{j}^{i}\in\mathcal{V}_{head}(u_{<j}^{i}),\forall u_{j}^{i}\in\mathrm{\textbf{u}}_{cont}^{i}.\end{split}$ (6)
@@
$\begin{split}&\mathcal{L}_{\mathrm{WAS}}(\mathrm{\textbf{u}}_{cont}^{i},\mathcal{O}^{i},\mathcal{U}^{i})\\ &=\sum_{z=1}^{s}{o_{z}^{i}\cdot f_{lm}{(\mathrm{\textbf{u}}_{cont}^{i}|\mathrm{\textbf{u}}_{z}^{i})}}\\ &=\prod_{j=1}^{q^{i}}{\text{WAS-score}(u_{j}^{i},u_{<j}^{i},\mathcal{O}^{i},\mathcal{U}^{i})},\end{split}$ (7)
@@
$\begin{split}&\text{WAS-score}(u_{j}^{i},u_{<j}^{i},\mathcal{O}^{i},\mathcal{U}^{i})\\ &=\begin{cases}\sum\limits_{z=1}^{s}{o_{z}^{i}\cdot f_{lm}(u_{j}^{i}|u_{<j}^{i},\mathrm{\textbf{u}}_{z}^{i})}&\text{if }u_{j}^{i}\in\mathcal{V}_{head}(u_{<j}^{i}),\\ -\inf&\text{otherwise}.\\ \end{cases}\end{split}$ (8)
@@ 4 Experiments
Method GSM8K HellaSwag
Retrieval (5-shot) 22.1 28.3
ParaICL (15-shot) 24.9 31.3
4.1 Datasets
4.2 Baselines
To provide a more comprehensive overview of where our framework stands, we use the following baselines: (1) Standard few-shot (Standard) (Brown et al., 2020): Directly generating the results based on few-shot demonstration examples. (2) Semantically sorted few-shot (Sorted) (Chen et al., 2023): Utilizing the same few-shot demonstration examples, this approach organizes the examples by the semantic similarity between each example’s question and the test question. Sorted+ indicates examples sorted in ascending similarity and Sorted- in descending similarity. (3) Coverage-based selection (CBS) (Gupta et al., 2023): CBS selects examples from more salient aspects such as reasoning patterns of the test input by using BERTScore-Recall. (4) Filter-then-search (LENS) (Li and Qiu, 2023b): LENS utilizes InfoScore to evaluate and progressively selects examples. (5) Parallel context window (PCW) (Ratner et al., 2023): Carving a long context into batches, PCW restricts the attention mechanism to apply within each batch. To ensure a fair comparison, we maintain the same number of batches as used in ParaICL. (6) Structured prompting (SP) (Hao et al., 2022b): In SP, demonstration examples are separately encoded with well-designed position embeddings and jointly attended by the test example using a rescaled attention mechanism.
4.3 Experiment Setup for ParaICL
When setting up experiments for ParaICL, we maintain consistency with the demonstration examples used in baseline methods to ensure a fair comparison. To demonstrate the effectiveness of ParaICL, we conduct experiments with varying numbers of total demonstration examples, specifically 3, 9, and 15, while keeping the count of parallel batches at three  [^2] ParaICL focuses on optimally using any number of available examples, and as such, we do not experiment with hundreds or thousands of examples.. Consequently, this results in 1, 3, and 5 demonstration examples per batch across the three settings. The impact of batch numbers will be elaborated in Section 5.1. Each experimental run is carried out with three random seeds for the selection of demonstration examples, and the results are averaged for reporting. Our main experiments are conducted using open-source models [^3] Due to cost concern, the validation of ParaICL’s effectiveness on closed-source models is conducted on a more compact dataset in Section 5.4., including Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2. We utilize supervised SimCSE (Gao et al., 2021) with BERT base to compute sentence embeddings.
@@ 4.4 Experiment Results
We present the experimental results of all datasets from Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 in Table 1. ParaICL consistently outperforms baseline methods on all datasets, demonstrating the effectiveness of our method. On reasoning tasks, the average improvements with Llama-2-7B-Chat are 1.2%, 2.7%, and 2.0% for 3-shot, 9-shot, and 15-shot settings, respectively. Mistral-7B-Instruct-v0.2 shows improvements of 1.0%, 2.9%, and 3.0% for the same settings, demonstrating ParaICL’s adaptability across different large language models. On commonsense NLI and coding tasks, ParaICL presents an average performance boost of 1.8% and 2.0% for Llama-2-7B-Chat, and 1.9% and 2.0% for Mistral-7B-Instruct-v0.2, respectively. Moreover, ParaICL consistently surpasses CBS and PCW in all settings and on virtually all datasets, underscoring the advantages of utilizing semantic information in our approach. An observation is made regarding the performance similarity between the Sorted+ and Sorted- methods. Sorted+ sometimes outperforms Sorted- and vice versa. This finding is consistent with prior research by Levy et al. (2023) and Liu et al. (2022). With ParaICL, this issue is lessened due to the reduced number of demonstration examples per batch.
@@
5 Ablation Studies and Analysis
5.1 Number of Batches
ParaICL keeps the context length for each batch manageable, allowing for the increasing of batch numbers as hardware capabilities allow. In this experiment, we set the number of demonstration examples in each batch as five, and progressively increase the number of batches. The results of Mistral-7B-Instruct-v0.2 on GSM8K is in Figure 4. With the increment in batch numbers, ParaICL’s improvements tend to converge at five batches. Further increase in batch numbers leads to instability in results. We attribute this to the higher likelihood of incorporating batches that negatively affect token selection. Consequently, we set the batch number to three in our main experiments to ensure stable performance improvements.